75600J of thermal energy is needed to raise the temperature of a 2.0kg block of ice removed from a freezer at−18∘ C to its melting point.
Calculate the specific heat capacity of ice suggested by these figures.
1
Expert's answer
2020-09-21T08:31:55-0400

Energy needed to change the temperature of the object of mass "m" with specific heat capacity "c" by "\Delta T" is "Q = c m \Delta T".

From the last formula, specific heat capacity is "c = \frac{Q}{m \Delta T} = \frac{75600 J}{2 kg \cdot 18^{\circ} C} = 2100 \frac{J}{kg ^{\circ}C}" .